During a recession, the price of goods and services goes down because of low demand. A company that makes Ethernet adapters iBull Built, a design/build engineering company that has always given year-end bonuses in the amount of $4,000 to each of itsCould you solve/fill the blank?
Answer #1

1)

Let us calculate the present worth of cost (PW1), if expansion is carried out after 1 year is given by

PW1=$110000/(1+11%)=$990990.99

Price if expansion is carried out now=PW=$830,000

Discount offered=990990.99.10-83000=$160990.99

2)

Next years' equivalent bonus =4000*(1+15%)+4000=$8600